# Xu Shen  and  Yuqiang Zheng Morningside Center of Mathematics Academy of Mathematics and Systems Science Chinese Academy of Sciences

### v2

@@ 3.3. Smoothness of $\zeta$
${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$${{\mathcal{G}^{\mathrm{spl}}_{k}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\operatorname{Spec}\,A}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{{\mathcal{G}_{0}^{\mathrm{spl}}}\text{-}\mathrm{Zip}^{\mu}_{\kappa}\otimes k}}$
The morphism $X\to\operatorname{Spec}\,A$ is a trivial $E$-torsor isomorphic to $\underline{I}_{ug}$ for an $A$-point $ug$ of $\mathcal{G}^{\mathrm{spl}}$. Let $U_{-}\subset\mathcal{G}^{\mathrm{spl}}_{k}$ be the opposite unipotent subgroup determined by $\mu$, then $A$ is isomorphic to the complete local ring of $U_{-}$ at identity. The trivialization induces an isomorphism and translates $\alpha$ into the morphism $\beta:\operatorname{Spec}\,A\times_{k}E_{k}\to\mathcal{G}^{\mathrm{spl}}_{k}$ given by for any $k$-scheme $T$, on $T$-points $\beta:(u,l,u_{+},u_{-})\mapsto lu_{+}ug(l^{(p)}u_{-})^{-1}$. Now the same method of the last paragraph of [67]*Theorem 4.1.2 shows the map on the tangent space at $x$ is surjective. ∎
@@ 3.4. Non-emptiness of EO strata
  • If $x$ is a minimal EO point of $\mathscr{A}^{\mathrm{spl}}(k)$, then $y$ is a minimal EKOR point of $\mathscr{A}(k)$.
@@
Both the involved $\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}}$-zip and $\operatorname{\mathcal{G}}_{0}^{\operatorname{rdt}}$-zips admit decompositions over the indexes $i$. We proceed according to the type of $i$. Case (AL/AU): For each $j$, write $\sum_{l}d_{i,j}^{l}=t_{j}d_{i}+s_{j},0\leq s_{j}<d_{i}$, and
@@
This forces $\omega_{x,i,j}\simeq\omega_{\mathrm{min}}\bigoplus\omega_{\mathrm{min}}^{\prime}$, so $y=\pi(x)$ is a minimal KR point, which is then a minimal EKOR point by the condition (1). Conversely, if $y$ is a minimal EKOR point, then by the above condition (1) and (2), there is a splitting structure $\underline{\operatorname{\mathscr{F}}_{\bullet}}$ of $y$, such that $x^{\prime}=(y,\underline{\operatorname{\mathscr{F}}_{\bullet}})$ minimal.
@@
By subsection A.5, the EKOR strata of $\mathscr{A}_{0}$ are non-empty. Hence, we have the existence of minimal EO points of $\operatorname{\mathscr{A}}_{k}^{\mathrm{spl}}$ by Proposition 3.11. The smoothness of $\zeta:\operatorname{\mathscr{A}}_{0}^{\mathrm{spl}}\to{\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}}}\text{-}\mathrm{Zip}^{\mu}_{\kappa}$ plus the fact that its image contains a minimal point imply the surjectivity of $\zeta$. ∎
@@ 3.5. The $\mu$-ordinary locus
Let $U\subset\mathscr{A}^{\mathrm{spl}}_{k}$ the maximal EO stratum, then by Theorem 3.13, $U$ is open dense in $\mathscr{A}^{\mathrm{spl}}_{k}$.
@@
The maximal KR stratum $\mathscr{A}_{k}^{w_{0}}$ can be described as follows: $x\in\mathscr{A}_{k}^{w_{0}}$ if and only if the following condition ( dependent on the type of the index $i$) holds:
@@
[…]
@@
By Proposition 3.15, the splitting structure over $U_{\rm KR}$ is unique, hence the EO strata in $\pi^{-1}(U_{\rm KR})$ is uniquely determined by the fiber of $\zeta_{1}$. In other words, the restriction of the universal $\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}}$-zip of type $\mu$ to $\pi^{-1}(U_{\rm KR})$ is uniquely determined by the associated $\operatorname{\mathcal{G}}_{0}^{\operatorname{rdt}}$-zip of type $\mu^{\prime}$. To the above diagram commutes, we need to show that this $\operatorname{\mathcal{G}}_{0}^{\operatorname{rdt}}$-zip is the pullback (under $\pi$) of the $\operatorname{\mathcal{G}}_{0}^{\operatorname{rdt}}$-zip on $U_{\rm KR}$ given by $\zeta_{2}=\zeta_{w_{0}}$.
@@
and $\Lambda_{k,i}$ decomposes as $W\oplus W^{\prime}$, with $W=k[\operatorname{\varepsilon}_{i}]^{\oplus d_{i}f_{i}}$, $W^{\prime}$ the dual of $W$. There are induced Frobenius morphism $F$ and Verschiebung morphism $V$ on $\Lambda_{k,i}$. The $F$-zip associated to $x$ (with $\operatorname{\mathcal{G}}_{i,0}^{\operatorname{rdt}}$-structure) induces an $F$-zip with $\kappa_{i}$-action $M=(W,C,D,\iota)$ given by
@@
The EO strata of $U_{\rm KR}$ is determined by an $F$-zip $(M^{\prime},C^{\prime},D^{\prime},\iota^{\prime})$ with $\operatorname{\mathcal{G}}_{i,0}^{\rm rdt}$-structure (recall that $\operatorname{\mathcal{G}}_{i,0}=\operatorname{Res}_{\kappa_{i}[\operatorname{\varepsilon}_{i}]|\mathbb{F}_{p}}H_{i},\operatorname{\mathcal{G}}_{i,0}^{\rm rdt}=\operatorname{Res}_{\kappa_{i}|\mathbb{F}_{p}}H_{i}$). More precisely, we have
@@
  • The morphism $F,V$ of $M$ restricts to the morphism
@@ 3.6. Hodge strata and pullbacks of Kottwitz-Rapoport strata
In [4], Bijakowski-Hernandez introduced a Hodge stratification. Here we give an interpretation of their Hodge stratification by our language.
Recall the local setup of [3, 4]. Let $L$ be a finite extension of $\operatorname{\mathbb{Q}}_{p}$ of degree $n=ef$, where $e$ is the ramification index, and $f$ the residue degree of $L|\operatorname{\mathbb{Q}}_{p}$. Let $\pi$ be a uniformizer of $\operatorname{\mathcal{O}}_{L}$. Let $G$ be a $p$-divisible group over $k$ with $\operatorname{\mathcal{O}}_{L}$ action $\iota$, such that its height is $nh$. This is the local geometric datum in the case of type (AL). In the case of type (C) or (AU), we require moreover there to be a polarization $\lambda:G\rightarrow G^{\vee}$ of $G$. Note that even in the type (AL) case, we have a natural polarization $\lambda:G\times G^{\vee}\rightarrow G^{\vee}\times G$. Let $(M,F)$ be the contravariant $F$-crystal associated to the $p$-divisible group $G$. Then $M$ is a finite free $W(k)$-module of rank $nh$ and $F:M\to M$ a $\sigma$-linear injective morphism. The $\operatorname{\mathcal{O}}_{L}$-action induces a decomposition of $W(k)$-module
@@
For each $0\leq i\leq h$, the $i$-th coordinate of Hodge polygon is:
@@
Note that our polygon is upper convex, where [3] uses lower convex polygon. All statements of loc. cit. about polygons still hold, after replacing minimal by maximal. The Hodge polygon of $\underline{G}=(G,\iota,\lambda)$ is defined as a polygon starting at $(0,0)$ and ending at $(h,d)$ such that for $0\leq i\leq h$
$\operatorname{Hdg}(\underline{G})(i)=\frac{1}{f}\sum_{\tau\in J}\operatorname{Hdg}_{\operatorname{\mathcal{O}}_{L},\tau}(M,F)(i),$
@@
$\omega_{G}\simeq M/FM$
as $\operatorname{\mathcal{O}}_{L}$-modules, the Hodge polygon of $\underline{G}$ depends only on the $\operatorname{\mathcal{O}}_{L}$-module $\omega_{G}$.
@@
Let $\underline{G}=(G,\iota,\lambda,\operatorname{Fil}^{\bullet})$ be a $p$-divisible group over a perfect field $k$ of characteristic $p$, with an $\operatorname{\mathcal{O}}_{L}$-action, plus a Pappas-Rapoport condition given by $\mu=(d_{\tau}^{l})_{\tau,l}$. We have three polygons associated to $\underline{G}$, the Newton polygon (upper convex) $\operatorname{Newt}(\underline{G})$ and the Hodge polygon $\operatorname{Hdg}(\underline{G})$ with respect to $\underline{G}$, and a PR polygon $\operatorname{PR}(\mu)$ determined by the Pappas-Rapoport condition with respect to $\mu$. One has the following inequalities:
$\operatorname{Newt}(\underline{G})\leq\operatorname{Hdg}(\underline{G})\leq\operatorname{PR}(\mu).$
Using our language, let $x\in\operatorname{\mathscr{A}}^{\mathrm{spl}}(k)$, and $G_{x}$ the $p$-divisible group associated to $x$. Then $G_{x}$ decomposes as direct sum $G_{x}=\bigoplus_{i=1}^{r}G_{x,i}$, and each $G_{x,i}$ is a $p$-divisible group over $k$ with $\operatorname{\mathcal{O}}_{F_{i}}$-action and polarization $\lambda$, plus a PR condition with respect to $\mu_{i}$. So for each $i$, there are three polygons $\operatorname{Newt}_{i}(x):=\operatorname{Newt}(G_{x,i}),\operatorname{Hdg}_{i}(x):=\operatorname{Hdg}(G_{x,i}),\operatorname{PR}(\mu_{i})$ associated to $x$. Note that for each $i$, we have an isomorphism of $\operatorname{\mathcal{O}}_{F_{i}}$-module $\omega_{x,i}\simeq\omega_{G_{x,i}}$.
@@
This follows from the description of maximal KR stratum in the proof of Proposition 3.15. ∎
@@
We give a group theoretic reformulation of Hodge polygon. For simplicity, we assume $r=1$ (the general case follows by combining the independent data for all $i\in I$) in our previous notations. Let $F_{1}$ be a finite extension over $\operatorname{\mathbb{Q}}_{p}$ of degree $n_{1}=e_{1}f_{1}$, with $e_{1}$ its ramification index and $f_{1}$ its residue degree. We also fix an unramified group $H$ over $F_{1}$ and let $G_{1}=\operatorname{Res}_{F_{1}|\operatorname{\mathbb{Q}}_{p}}H$. The maximal parahoric subgroup of $G_{1}(\mathbb{Q}_{p})$ is $K=H(\operatorname{\mathcal{O}}_{F_{1}})$, where we use the same $H$ as the reductive model of $H$. Let ˘$\breve{\operatorname{\mathbb{Q}}}_{p}$ be the completion of maximal unramified extension of ${\operatorname{\mathbb{Q}}}_{p}$. For simplicity, we will write $\breve{G}_{1}=G_{1}(\breve{\operatorname{\mathbb{Q}}}_{p})$ and similar symbols for other groups.
@@
where $\Gamma_{0}=\operatorname{Gal}(\overline{\operatorname{\mathbb{Q}}}_{p}/\breve{\operatorname{\mathbb{Q}}}_{p})$ and $W_{0}$ the relative Weyl group of $G_{\breve{\operatorname{\mathbb{Q}}}_{p}}$. Let $\breve{K}\subset\breve{G}_{1}$ be the induced parahoric subgroup and
@@
Recall the $\mu$-admissible set $\operatorname{Adm}(\mu)_{K}\subset W_{K}\backslash\widetilde{W}/W_{K}$. Restricting $\operatorname{Hdg}$ to it gives the following proposition.
@@
Let $(M,F)$ be the $F$-crystal with additional structure attached to $x$, which depends only on $\pi(x)$. The isomorphism class of $M/FM$ as $\operatorname{\mathcal{O}}_{L}\otimes W(k)$-module is given by an element in $\operatorname{Adm}(\mu)_{K}$, consisting of a tuple in $X_{*}(T)_{\Gamma_{0}}$. The average of the tuple in $X_{*}(T)_{\Gamma_{0}}$ gives the Hodge polygon of $F$-crystal $(M,F)$, which is (as usual) viewed as an element of $X_{*}(T_{1})_{\Gamma_{0},\operatorname{\mathbb{Q}}}^{+}$. ∎
@@
Recall that Corollary 3.19 says that the maximal Hodge stratum is exactly the preimage of the maximal Kottwitz-Rapoport stratum. One checks easily that $\operatorname{Hdg}(w_{0})=\operatorname{PR}(\mu)$.

### v3

@@ 3.3. Smoothness of $\zeta$
[…]
The morphism $X\to\operatorname{Spec}\,A$ is a trivial $E$-torsor isomorphic to $\underline{I}_{ug}$ for an $A$-point ugug of $\mathcal{G}^{\mathrm{spl}}$. Let $U_{-}\subset\mathcal{G}^{\mathrm{spl}}_{k}$ be the opposite unipotent subgroup determined by $\mu$, then $A$ is isomorphic to the complete local ring of $U_{-}$ at identity. The trivialization induces an isomorphism and translates $\alpha$ into the morphism $\beta:\operatorname{Spec}\,A\times_{k}E_{k}\to\mathcal{G}^{\mathrm{spl}}_{k}$ given by for any $k$-scheme $T$, on $T$-points $\beta:(u,l,u_{+},u_{-})\mapsto lu_{+}ug(l^{(p)}u_{-})^{-1}$. Now the same method of the last paragraph of [70]*Theorem 4.1.2 shows the map on the tangent space at $x$ is surjective. ∎
@@ 3.4. Non-emptiness of EO strata
  • If $x$ is a minimal EO point of $\mathscr{A}^{\mathrm{spl}}(k)$ (i.e. $x$ lies in the minimal EO stratum), then $y$ is a minimal EKOR point of $\mathscr{A}(k)$ (i.e. $y$ lies in the minimal EKOR stratum, cf. [59] 1.2.5).
@@
Both the involved $\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}}$-zip and $\operatorname{\mathcal{G}}_{0}^{\operatorname{rdt}}$-zips admit decompositions over the indices $i$. We proceed according to the type of $i$.
Case (AL/AU): For each $j$, write $\sum_{l}d_{i,j}^{l}=t_{j}d_{i}+s_{j}$, where $0\leq s_{j}<d_{i}$, and
@@
This forces
$\omega_{x,i,j}\simeq\omega_{\mathrm{min}}\bigoplus\omega_{\mathrm{min}}^{\prime},$
so $y=\pi(x)$ is a minimal KR point, which is then a minimal EKOR point by the condition (1). Conversely, if $y$ is a minimal EKOR point, then by the above condition (1) and (2), there is a splitting structure $\underline{\operatorname{\mathscr{F}}_{\bullet}}$ of $y$, such that $x^{\prime}=(y,\underline{\operatorname{\mathscr{F}}_{\bullet}})$ minimal.
@@
By Proposition A.15, the EKOR strata of $\mathscr{A}_{0}$ are non-empty. Hence, minimal EO points of $\operatorname{\mathscr{A}}_{k}^{\mathrm{spl}}$ exist by Proposition 3.11. The smoothness of $\zeta:\operatorname{\mathscr{A}}_{0}^{\mathrm{spl}}\to{\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}}}\text{-}\mathrm{Zip}^{\mu}_{\kappa}$ plus the fact that its image contains a minimal point imply the surjectivity of $\zeta$. ∎
@@ 3.5. The $\mu$-ordinary locus
Let $U\subset\mathscr{A}^{\mathrm{spl}}_{k}$ be the maximal EO stratum, then by Theorem 3.13, $U$ is open dense in $\mathscr{A}^{\mathrm{spl}}_{k}$.
@@
The maximal KR stratum $\mathscr{A}_{k}^{w_{0}}$ can be described as follows: $x\in\mathscr{A}_{k}^{w_{0}}(k)$ if and only if the following condition ( dependent on the type of the index $i$) holds:
@@
${\pi^{-1}(U_{\rm KR})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}\scriptstyle{\simeq}$$\scriptstyle{\zeta_{1}}$${{\operatorname{\mathcal{G}}_{0}^{\operatorname{rdt}}}\text{-}\mathrm{Zip}^{\mu^{\prime}}_{k}}$${U_{\rm KR}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\zeta_{2}}$
@@
By Proposition 3.15, the splitting structure over $U_{\rm KR}$ is unique, hence the EO strata in $\pi^{-1}(U_{\rm KR})$ are uniquely determined by the fiber of $\zeta_{1}$. In other words, the restriction of the universal $\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}}$-zip of type $\mu$ to $\pi^{-1}(U_{\rm KR})$ is uniquely determined by the associated $\operatorname{\mathcal{G}}_{0}^{\operatorname{rdt}}$-zip of type $\mu^{\prime}$. To prove that the above diagram commutes, we need to show that this $\operatorname{\mathcal{G}}_{0}^{\operatorname{rdt}}$-zip is the pullback (under $\pi$) of the $\operatorname{\mathcal{G}}_{0}^{\operatorname{rdt}}$-zip on $U_{\rm KR}$ given by $\zeta_{2}=\zeta_{w_{0}}$.
@@
and $\Lambda_{k,i}$ decomposes as $W\oplus W^{\prime}$, with $k[\operatorname{\varepsilon}_{i}]^{\oplus d_{i}f_{i}}, WW^{\prime}$ the dual of $W$. There are induced Frobenius morphism $F$ and Verschiebung morphism $V$ on $\Lambda_{k,i}$. The $F$-zip associated to $x$ (with $\operatorname{\mathcal{G}}_{i,0}^{\operatorname{rdt}}$-structure) induces an $F$-zip with $\kappa_{i}$-action $M=(W,C,D,\iota)$ given by
@@
The EO strata inside $U_{\rm KR}$ are determined by an $F$-zip $(M^{\prime},C^{\prime},D^{\prime},\iota^{\prime})$ with $\operatorname{\mathcal{G}}_{i,0}^{\rm rdt}$-structure (recall that $\operatorname{\mathcal{G}}_{i,0}=\operatorname{Res}_{\kappa_{i}[\operatorname{\varepsilon}_{i}]|\mathbb{F}_{p}}H_{i},\operatorname{\mathcal{G}}_{i,0}^{\rm rdt}=\operatorname{Res}_{\kappa_{i}|\mathbb{F}_{p}}H_{i}$). More precisely, we have
@@
  • The morphisms $F,V$ of $M$ restrict to the morphisms
@@ 3.6. Hodge strata and pullbacks of Kottwitz-Rapoport strata
In [6], Bijakowski-Hernandez introduced a Hodge stratification. Here, we provide an interpretation of their Hodge stratification in our language.
Recall the local setup of [5, 6]. Let $L$ be a finite extension of $\operatorname{\mathbb{Q}}_{p}$ of degree $n=ef$, where $e$ is the ramification index, and $f$ the residue degree of $L|\operatorname{\mathbb{Q}}_{p}$. Let $\pi$ be a uniformizer of $\operatorname{\mathcal{O}}_{L}$. Let $\mathbf{G}$ be a $p$-divisible group over $k$ with $\operatorname{\mathcal{O}}_{L}$ action $\iota$, such that its height is $nh$. This is the local geometric datum in the case of type (AL). In the case of type (C) or (AU), we require moreover there to be a polarization $\lambda:\mathbf{G}\rightarrow\mathbf{G}^{\vee}$ of $\mathbf{G}$. Note that even in the type (AL) case, we have a natural polarization $\lambda:\mathbf{G}\times\mathbf{G}^{\vee}\rightarrow\mathbf{G}^{\vee}\times\mathbf{G}$. Let $(M,F)$ be the contravariant $F$-crystal associated to the $p$-divisible group $\mathbf{G}$. Then $M$ is a finite free $W(k)$-module of rank $nh$ and $F:M\to M$ a $\sigma$-linear injective morphism. The $\operatorname{\mathcal{O}}_{L}$-action induces a decomposition of $W(k)$-module
@@
For each $0\leq i\leq h$, the $i$-th coordinate of the Hodge polygon is:
@@
Note that our polygon is upper convex, whereas [5] uses lower convex polygon. All statements of loc. cit. about polygons still hold, after replacing minimal by maximal. The Hodge polygon of $\underline{\mathbf{G}}=(\mathbf{G},\iota,\lambda)$ is defined as a polygon starting at $(0,0)$ and ending at $(h,d)$ such that for $0\leq i\leq h$
$\operatorname{Hdg}(\underline{\mathbf{G}})(i)=\frac{1}{f}\sum_{\tau\in J}\operatorname{Hdg}_{\operatorname{\mathcal{O}}_{L},\tau}(M,F)(i),$
@@
$\omega_{\mathbf{G}}\simeq M/FM$
as $\operatorname{\mathcal{O}}_{L}$-modules, the Hodge polygon of $\underline{\mathbf{G}}$ depends only on the $\operatorname{\mathcal{O}}_{L}$-module $\omega_{\mathbf{G}}$.
@@
Let $\underline{\mathbf{G}}=(\mathbf{G},\iota,\lambda,\operatorname{Fil}^{\bullet})$ be a $p$-divisible group over a perfect field $k$ of characteristic $p$, with an $\operatorname{\mathcal{O}}_{L}$-action, plus a Pappas-Rapoport condition given by $\mu=(d_{\tau}^{l})_{\tau,l}$. We have three polygons associated to $\underline{\mathbf{G}}$, the Newton polygon (upper convex) $\operatorname{Newt}(\underline{\mathbf{G}})$ and the Hodge polygon $\operatorname{Hdg}(\underline{\mathbf{G}})$ with respect to $\underline{\mathbf{G}}$, and a PR polygon $\operatorname{PR}(\mu)$ determined by the Pappas-Rapoport condition with respect to $\mu$. One has the following inequalities:
$\operatorname{Newt}(\underline{\mathbf{G}})\leq\operatorname{Hdg}(\underline{\mathbf{G}})\leq\operatorname{PR}(\mu).$
Using our language, let $x\in\operatorname{\mathscr{A}}^{\mathrm{spl}}(k)$, and $\mathbf{G}_{x}$ the $p$-divisible group associated to $x$. Then $\mathbf{G}_{x}$ decomposes as direct sum $\mathbf{G}_{x}=\bigoplus_{i=1}^{r}\mathbf{G}_{x,i}$, and each $\mathbf{G}_{x,i}$ is a $p$-divisible group over $k$ with $\operatorname{\mathcal{O}}_{F_{i}}$-action and polarization $\lambda$, plus a PR condition with respect to $\mu_{i}$. So for each $i$, there are three polygons $\operatorname{Newt}_{i}(x):=\operatorname{Newt}(\mathbf{G}_{x,i}),\operatorname{Hdg}_{i}(x):=\operatorname{Hdg}(\mathbf{G}_{x,i}),\operatorname{PR}(\mu_{i})$ associated to $x$. Note that for each $i$, we have an isomorphism of $\operatorname{\mathcal{O}}_{F_{i}}$-module $\omega_{x,i}\simeq\omega_{\mathbf{G}_{x,i}}$.
@@
This follows from the description of the maximal KR stratum in the proof of Proposition 3.15. ∎
@@
We give a group theoretic reformulation of Hodge polygon. For simplicity, we assume $r=1$ (the general case follows by combining the independent data for all $i\in I$) in our previous notations. Let $F_{1}$ be a finite extension over $\operatorname{\mathbb{Q}}_{p}$ of degree $n_{1}=e_{1}f_{1}$, with $e_{1}$ its ramification index and $f_{1}$ its residue degree. We also fix an unramified group $H$ over $F_{1}$ and let $G_{1}=\operatorname{Res}_{F_{1}|\operatorname{\mathbb{Q}}_{p}}H$. The maximal parahoric subgroup of $G_{1}(\mathbb{Q}_{p})$ is $K=H(\operatorname{\mathcal{O}}_{F_{1}})$, where we use the same $H$ as the reductive model of $H$. Let $\breve{\operatorname{\mathbb{Q}}}_{p}$ be the completion of maximal unramified extension of ${\operatorname{\mathbb{Q}}}_{p}$. For simplicity, we will write $\breve{G}_{1}=G_{1}(\breve{\operatorname{\mathbb{Q}}}_{p})$ and similar symbols for other groups.
@@
where $\Gamma_{0}=\operatorname{Gal}(\overline{\operatorname{\mathbb{Q}}}_{p}/\breve{\operatorname{\mathbb{Q}}}_{p})$ and $W_{0}$ is the relative Weyl group of $G_{\breve{\operatorname{\mathbb{Q}}}_{p}}$. Let $\breve{K}\subset\breve{G}_{1}$ be the induced parahoric subgroup and
@@
[…]
@@
Let $(M,F)$ be the $F$-crystal with additional structure attached to $x$, which depends only on $\pi(x)$. The isomorphism class of $M/FM$ as $\operatorname{\mathcal{O}}_{L}\otimes W(k)$-module is given by an element in $\operatorname{Adm}(\mu)_{K}$, consisting of a tuple in $X_{*}(T)_{\Gamma_{0}}$. The average of the tuple in $X_{*}(T)_{\Gamma_{0}}$ gives the Hodge polygon of the $F$-crystal $(M,F)$, which is (as usual) viewed as an element of $X_{*}(T_{1})_{\Gamma_{0},\operatorname{\mathbb{Q}}}^{+}$. ∎
@@
Recall that Corollary 3.19 states that the maximal Hodge stratum is exactly the preimage of the maximal Kottwitz-Rapoport stratum. One checks easily that $\operatorname{Hdg}(w_{0})=\operatorname{PR}(\mu)$.